OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \normalsize \text{prove that the equation}\hspace{.33em}\\~\\ \dfrac{1}{x}+\dfrac{1}{y}=2\hspace{.33em}\\~\\ \normalsize \text{has no integer solution ,except }\hspace{.33em}\\~\\ x=y=1\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

this equation can be written in the form : x= y/(2y-1) now you can see that 2y-1 is bigger than y for all integers values but 1. hence x wont be an integer for y integer other than 1.

OpenStudy (anonymous):

note that the same thing can be said for negative y. and remember x cant be 0 and also y cant be 0.

Parth (parthkohli):

We can be confident that no other solutions can be found when both have the same sign, because a fractional value is always < 1 for positive integers and in the case of negative, it's pretty obvious how the left-hand-side is negative unlike the right side. Let's say that x and y are of opposite signs where \(x>0, y<0\) WLOG.\[\dfrac{1}{x} = 2 - \frac{1}{y}\]LHS < 1 and RHS > 2. Hence proved.

Parth (parthkohli):

Case 1: x > 0, y > 0. LHS \(\le 2\). Case 2: x < 0, y < 0. LHS \(< 0\). Case 3: x < 0, y > 0. As shown above.

OpenStudy (anonymous):

1/x+1/y=2 multiply both sides by xy y+x=2xy add negative y for both sides x=2xy-y x=y(2x-1) y=x/(2x-1)

OpenStudy (michele_laino):

we can perform a demonstration by absurd, as below: let's suppose that x is an integer m which is greater than 1, namely x=m>1 and y is another integer equal to 1, namely y=1. Now we have: \[\frac{ 1 }{ m }<1\] on the other hand if I substitute those above quantities into the starting expression, I get: \[\frac{ 1 }{ m }+1=2\] or: \[\frac{ 1 }{ m }=1\] which is a contraddiction.

OpenStudy (mathmath333):

thanks all